OpenStudy (anonymous):

How do you simplify imaginary numbers? i^93

OpenStudy (shubhamsrg):

make use of this : i^2 = -1 i^4 =1 i^3 =-1

OpenStudy (lgbasallote):

have you tried doing \[\LARGE i^{92} \times i?\]

OpenStudy (anonymous):

is i to any even exponent = -1?

OpenStudy (lgbasallote):

\[\huge i^{92} \times i \implies (i^2)^{46} \times i \implies (-1)^{46} \times i = 1 \times i\]

OpenStudy (lgbasallote):

so to your question...no i^2 is -1...you have to simplify it further

OpenStudy (lgbasallote):

do you get the logic?

OpenStudy (anonymous):

hmm, not bad

OpenStudy (lgbasallote):

i think i raised to any even exponent other than 2 becomes 1...i raised to odd number other than 3 results in just i i think...im not sure...it's better to just do it slowly to be sure

OpenStudy (anonymous):

k thnx bye